# Deep Learning assisted Port-Cycling based Channel Sounding for Precoder Estimation in Massive MIMO Arrays

### v1

@@ IV-A Link-Level Simulation Assumptions
The simulator is configured to operate on 3GPP Urban Macro (UMa) Line-of-Sight (LoS) scenario with an inter-site distance of 200m [9]. It is a single cell setup, sectorized into three sectors. The BS is equipped with 128 antennas in a uniform planar array structure while the UE is configured with 4 antennas. The BS antenna array is partitioned into sub-panel based sub-arrays with spatial domain reduction factor ($\rho=4$). The periodicity ($P$) of downlink CSI-RS is set to $20ms$. In this simulation, we consider the rank of the channel be to one i.e., $\nu=1$. The key simulation parameters are summarized in Table II.
@@ IV-B CsiAdaNet Model
The dataset consists of $30,000$ samples and is split into training, validation, and test sets in a $70:20:10$ ratio. The model is trained under multiple noise conditions by uniformly applying SNR levels from $\{-5,0,5,10,15\}$dB to the training data to increase its robustness. During testing, the model is evaluated independently at each SNR level in $\{-20,-15,-10,-5,0,5,10,15,20\}$ dB to assess performance across a wide SNR range. A batch size of $256$ is considered for $250$ epochs using the Adam optimizer with a learning rate of $10^{-3}$. The CsiAdaNet model is trained using a weighted multi-task loss, with a loss weights ($\lambda_{\mathcal{B}}:0.75,\lambda_{\mathcal{I}}:1.5,\;\lambda_{\mathcal{P}}:0.75,\;\lambda_{\mathcal{\varphi}}:1.25)$. A higher weightage is assigned to beam indices ($\lambda_{\mathcal{I}}$) to emphasize beam selection, as any error in that propagates directly in precoder construction. All of the prediction heads employ sparse categorical cross-entropy losses with sparse categorical accuracy as the evaluation metric. Early stopping is deployed with a patience of 25 and adaptive learning rate reduction are used to stabilize training and avoid overfitting.
To perform prediction, each sub-array is sounded at a different time instant, covering the full array over $\rho$ occasions. During training, all possible sounding orders are to be considered to avoid bais towards a specific sequence; For $\rho=4$ this results in $4!=24$ permutations. The port cycled estimates collected over $(\rho\times P)\;ms$, forms the input to the DL model, ensuring generalization against sounding order variations.
@@ IV-C Metrics
Squared Generalized Cosine Similarity(SGCS): It captures the directional alignment between the Type-II precoder traditional or CsiAdaNet predicted, and the dominant eigen vector of the corresponding covariance matrix with the actual number of ports of the channel. It ranges between [0,1].
@@
Beamforming Gain (BF): It measures the effective power captured by the Type-II precoder traditional or CsiAdaNet based, relative to the optimal eigen-beamformer.
@@ IV-D1 Precoder Alignment and Beamforming Performance
Figs. 5(a) and 5(b) evaluate the alignment and beamforming performance of the CsiAdaNet predicted precoder. As shown in Fig. 5(a), the DL based approach closely follows the traditional technique in terms of SGCS across all SNR regimes. At low SNR, the SGCS is lower (i.e., <0.2<0.2) due to noise-dominated sub-array CSI, but steadily improves with SNR and saturates at around $0.7$ in medium-to-high SNR regime, approaching the baseline trend. Consistently, Fig. 5(b) shows that beamforming gain achieved using CsiAdaNet predicted precoder closely matches the traditional baseline, with a small gap at high SNR (within $\approx 0.03$). The drop at low SNR arises from the higher sensitivity of beamforming gain to angular and amplitude estimation errors.
@@ IV-D2 Analysis of CSI feedback quantities
Figs. 5(c)–5(f) summarize the prediction accuracy of the individual CSI feedback quantities estimated by CSIAdaNet. In Fig. 5(c), beamset prediction is highly reliable even at moderate SNR levels, which can be attributed to the fact that its a coarse spatial resolution and has inherent robustness to noise and spatial sparsity. This supports the assumption that dominant spatial characteristics remain stable over the port-cycling duration. In Fig. 5(d), the beam indices prediction improves more gradually with SNR, reflecting the increased sensitivity associated with finer angular resolution. The wideband amplitude accuracy in Fig. 5(e) follows a similar trend, as amplitude estimation depends on the correctness of the selected beam indices. Finally, Fig. 5(f) indicates that wideband phase prediction is the a challenging task, particularly at low SNR, due to its reliance on coherent combining. Nevertheless, accurate phase estimation is achieved at higher SNR levels, demonstrating the effectiveness of temporally aggregated port-cycled CSI measurements.
@@ IV-D3 Complexity Analysis
This subsection compares the computational complexity of the proposed CSIAdaNet model with the conventional covariance based methods. Conventional approaches compute eigenvalue decomposition (EVD) of the full covariance matrix $R_{N_{t}}\in\mathbb{C}^{N_{t}\times N_{t}}$, incurring a complexity of $O(N^{3}_{t})$, along with an additional, $O(O_{1}O_{2}N^{2}_{t})$ operations for selecting beam related metrics. In contrast, the proposed DL based approach performs inference on sub-panel observations, that scales as a $O(\rho D^{2})$, where $\rho$ denotes the number of subpanels and $D$ is the sub-panel feature dimension. The proposed approach remains independent of the full array size $N_{t}$, providing improved scalability for large-array systems.
@@ V Conclusion
This work presents a DL-assisted port-cycling based precoder estimation framework that jointly exploits spatial diversity across antennas and temporal diversity enabled by channel stationarity to reduce instantaneous CSI transmission overhead in large MIMO systems. Rather than sounding all ports simultaneously, spatially disjoint sub-panels are activated sequentially, and the resulting partial CSI measurements are coherently aggregated over time. The proposed CsiAdaNet model leverages the resulting spatial–temporal correlations to reconstruct Type-II codebook quantities from reduced CSI observations. Link-level simulations conducted in a 3GPP-compliant setup demonstrate that the proposed approach achieves near-optimal beam alignment and beamforming gain at moderate to high SNR levels, with only marginal degradation under noise-limited regions. Although all antenna ports are eventually sounded across multiple occasions, the number of active ports at any given instant is reduced by a factor of $\rho$, leading to reduced instantaneous overhead and RF chain utilization. This distinction between instantaneous and aggregate CSI overhead highlights a key design consideration for next-generation large-scale MIMO systems, and demonstrates the effectiveness of exploiting spatial and temporal diversity for scalable and energy-efficient CSI acquisition.

### v2

@@ IV-A Link-Level Simulation Assumptions
The simulation is conducted using a 3GPP compliant link-level simulator under an Urban Macro (UMa) line-of-sight scenario with an 200m inter-site distance [9]. A single-cell, three-sector deployment is considered, with a 128-antenna uniform planar array at the BS and four antennas at the UE. The BS array is partitioned into sub-panels with spatial reduction factor ($\rho=4$) and the CSI-RS periodicity is set to 20 ms, and single layer transmission is assumed, i.e ($\nu=1$). The key simulation parameters are summarized in Table II.
@@ IV-B CsiAdaNet Model
The dataset consists of $30,000$ samples, split into training, validation, and test sets in a $70:20:10$ ratio. To improve robustness, training is performed under multiple SNR conditions uniformly sampled from $\{-5,0,5,10,15\}$ dB, while evaluation is conducted over a wider SNR range of 20-20 to $20$ dB. The model is trained for $250$ epochs using Adam optimizer with a learning rate of $10^{-3}$ and a batch size of $256$ employing a weighted multi-task loss, with a loss weights ($\lambda_{\mathcal{B}}:0.75,\lambda_{\mathcal{I}}:1.5,\;\lambda_{\mathcal{P}}:0.75,\;\lambda_{\mathcal{\varphi}}:1.25)$. A higher weightage is assigned to beam indices ($\lambda_{\mathcal{I}}$) to emphasize any error in that propagates directly in precoder construction. All heads use sparse categorical cross-entropy with accuracy as the evaluation metric, while early stopping and adaptive learning-rate reduction are applied to prevent overfitting. Sub-arrays are sequentially sounded for port cycling, with all ($\rho!$) sounding permutations used during training to ensure model robustness.
@@ IV-C Metrics
Squared Generalized Cosine Similarity(SGCS): measures the directional alignment between the traditional or CsiAdaNet predicted Type-II precoder and the dominant eigen vector of the channel, with values ranging from 0 to 1.
@@
Beamforming Gain (BF): measures the effective power captured by the traditional or CsiAdaNet based Type-II precoder, relative to the optimal eigen-beamformer.
@@ IV-D1 Precoder Alignment and Beamforming Performance
Figs. 5(a) and 5(b) evaluate the alignment and beamforming performance of CsiAdaNet. As shown in Fig. 5(a), the DL based approach closely follows the conventional baseline across all SNRs, with SGCS increasing from below $0.2$ at low SNR to approximately $0.7$ in the medium-to-high SNR region. Consistently, Fig. 5(b) shows that beamforming gain achieved using CsiAdaNet, exhibiting close matches to the traditional baseline, with a small gap at high SNR (within $\approx 0.03$). The drop at low SNR is primarily due to noise-sensitive estimation.
@@ IV-D2 Analysis of CSI feedback quantities
Figs. 5(c)–5(f) summarize the prediction accuracy of the individual CSI feedback quantities. In Fig. 5(c), beamset prediction remains highly reliable even at moderate SNR levels, due to its coarse spatial resolution and inherent robustness to noise, validating the port-cycling assumption. In Fig. 5(d), the beam indices accuracy improves more gradually with SNR, reflecting the finer angular sensitivity. The wideband amplitude accuracy in Fig. 5(e) follows a similar trend, as amplitude estimation depends on the correctness of the beam selection. Finally, Fig. 5(f) indicates that wideband phase prediction is more challenging at low SNR but achieves high accuracy at moderate-to-high SNR, demonstrating the effectiveness of temporally aggregated port-cycled CSI.
@@ IV-D3 Complexity Analysis
This subsection compares the computational complexity of the CsiAdaNet model with conventional methods. Conventional approaches require eigenvalue decomposition (EVD) of the full covariance matrix $R\in\mathbb{C}^{N_{t}\times N_{t}}$, incurring $O(N^{3}_{t})$ complexity. In contrast, CsiAdaNet performs inference on sub-panel observations, with complexity $O(\rho D^{2})$, independent of the full array size $N_{t}$, providing improved scalability for large-array systems.
@@ V Conclusion
This work presents a DL-assisted port-cycling based precoder estimation that exploits both spatial diversity across antennas and temporal diversity arising from port-cycled measurements to reduce instantaneous CSI overhead in large MIMO systems. Rather than sounding all antenna ports simultaneously, spatially disjoint sub-panels are activated sequentially, and the partial CSI measurements are temporally aggregated . The proposed CsiAdaNet model leverages these spatial–temporal correlations to reconstruct Type-II codebook quantities from reduced CSI observations. Link-level simulations in a 3GPP-compliant setup demonstrate near-optimal beam alignment and beamforming gain at moderate to high SNR levels, with only marginal degradation under noisy regions. Although all antenna ports are eventually sounded across multiple occasions, the number of active ports at any given instant is reduced by a factor of $\rho$, reducing instantaneous overhead and RF chain usage, highlighting a key design consideration for scalable and energy-efficient CSI acquisition in next-generation large-scale MIMO systems.